Flamo:

The graph of y = -3x + 2 intersects the graph of an exponential function at (-1, 5). Which of the following could be the exponential function?

Flamo:

I'm Stuck...

Mehek:

What are the options given

Flamo:

Umm

Flamo:

y = 10(2)^2 y = 10(0.5)^2 y = 5^x y = 5(2)^x

Flamo:

Those are the answer Choices.

Flamo:

But I need to know how to do it...

bm717:

\[y=10(2)^2 , y=10(0.5)^2, y=5^x, y=5(2)^x\] Just to make it easier to see.

Mehek:

If it intersects at (-1, 5), that means that point is on both lines

Flamo:

o

Mehek:

So you can graph y = -3x + 2 and the options and see which one goes on (-1, 5) with y = -3x + 2 Or you can substitute -1 for x and 5 for y in the options and if both sides are equal it means the point is a solution of the exponential function, making it your answer

Flamo:

Okay. Thanks!